OpenStudy (anonymous):

p=a+b+c;p=41,a=11,b=12 c = ? perimeter of a triangle ?

jimthompson5910 (jim_thompson5910):

it asks "perimeter of a triangle ?" but p = 41 is given

jimthompson5910 (jim_thompson5910):

so that seems a bit odd

jimthompson5910 (jim_thompson5910):

anyways, it looks like they just want the value of c p = a+b+c 41 = 11+12+c solve for c to get c = ???

OpenStudy (anonymous):

im confused

jimthompson5910 (jim_thompson5910):

where at?

OpenStudy (anonymous):

all of it! lol

jimthompson5910 (jim_thompson5910):

do you see how I went from p = a+b+c to 41 = 11+12+c ?

jimthompson5910 (jim_thompson5910):

If not, then all I did was plug in p = 41, a = 11, b = 12

jimthompson5910 (jim_thompson5910):

After you plug those values in, you have to isolate c

jimthompson5910 (jim_thompson5910):

what do you get when you do that?

OpenStudy (anonymous):

18

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

as a check, p = a+b+c 41 = 11+12+18 41 = 41 so the solution is confirmed